What are all the zeros of the function:f(x)=4(-4x+6)(6x+4)(6x-8)?
Answer by stanbon(48558)

What are all the zeros of the function:f(x)=4(-4x+6)(6x+4)(6x-8)?
-------
Set each factor = to zero and solve for x:
Ex: If -4x+6=0 then 4x=6 and x=3/2
Ex: If 6x+4=0 then 6x=-4 and x= -2/3
====================
Cheers,
Stan H.